OpenStudy (anonymous):

for any time t≥0, the acceleration of a particle moving on the x-axis is a(t)=t+sin(t). At t=0. the velocity of the particle is -2. For what value t will the velocity of the particle be zero?

OpenStudy (anonymous):

\[a(t) = \frac{ dv }{ dt }\] thus:\[\frac{ dv }{ dt } = t + \sin t \] \[dv = tdt + \sin(t) dt\] Now integrate both sides

OpenStudy (anonymous):

\[\int\limits{}dv = \int\limits{}tdt + \int\limits{}\sin(t)dt\]

OpenStudy (anonymous):

to get velocity right